EVA Report – February 1st

EVA Report:
SOL: 03
Date: 02/01/2017
Person filling out the report: Sai Arun Dharmik
Crew members involved in the EVA: Sai Arun Dharmik and Saroj Kumar
EVA leader: Sai Arun Dharmik
Begin: 11:45 Hrs
End: 15:15 Hrs
Type of EVA: Deimos + 1 ATV (300) + Walking
Purpose: Collection of rock samples to be measured magnetic susceptibility
UTM Coordinates: Northing: 4255600 , Easting : 518400
Summary: We started the EVA at 11:45 am. The temperature was around 12
degree Celsius outside the Hab and the sky was clear. We have
collected few samples from the URC North site area on our route to
Lith Canyon. The samples collected at URC North site were primarily
transported due to weathering, hence we cannot attribute them to that
location. Later we collected few samples of sandstones and shales near
the Lith canyon region. The samples will be analysed tomorrow
(02.02.2017) and a detailed report will be sent.